2)Maui Macadamia Inc. has a monopoly in the macadamia nut industry. The demand curve, marginal revenue and marginal cost curve for macadamia nuts are given as follows: P = 360 - 4Q. MR = 360 - 8Q. MC = 4Q

e) At the profit maximizing level of output, what is the deadweight loss?

f) What is the maximum amount that Maui Macadamia would be willing to spend in order to maintain its monopoly through rent seeking?
Answer #1

③ P-360-40 MR=360-80 Mc-40 Profit is max when mean 360-89 ud 0 = 360 30 & brice-360-4130) bin - 240 4 Mchto Unele perfect com